OpenStudy (anonymous):

write 3 equivalent fractions 3/5 -4/4 -2/3 4/7

OpenStudy (shinalcantara):

say for 3/5, all you have to do is to multiply both the numerator and the denominator with a number equal to 1 let's say \[\frac{ 3 }{ 5 } \times \frac{ 4 }{ 4 } = \frac{ 12 }{ 20 }\] 12/20 is one of 2/3's equivalent fractions

OpenStudy (shinalcantara):

do the same and you'll have the answers